OpenStudy (anonymous):

y-5=3(x+1) in slope intercept form

OpenStudy (anonymous):

that is slope intercept isn't it

OpenStudy (anonymous):

no its point slope

pooja195 (pooja195):

Hi welcome to openstudy :) Start by distributing 3 to the x+1 3(x) =? 3(1)=?

OpenStudy (anonymous):

@pooja195 3

OpenStudy (anonymous):

@pooja195 and 3x

pooja195 (pooja195):

well done y-5+5=3x+3+5 Now just add 5 to both sides :) What are you left with?

OpenStudy (anonymous):

y=3x+8? @pooja195

pooja195 (pooja195):

Yep good job :)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

thanks :)

pooja195 (pooja195):

yw ^-^

OpenStudy (anonymous):

i have one more question @pooja195

pooja195 (pooja195):

Post it :)

OpenStudy (anonymous):

which of the following is a correct equation for the line passing through (-3,2) slope m=2/3 A. Y-2=2/3(X+3) B. 2X-3Y=-12 C. Y= 2/3X D. Y=2/3X+4 Choose all that apply @pooja195

pooja195 (pooja195):

@Mehek14

Mehek (mehek14):

which ones do you think?

OpenStudy (anonymous):

d and a

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

A is correct

OpenStudy (anonymous):

thats it?

Mehek (mehek14):

hold on

Mehek (mehek14):

B works as well

Mehek (mehek14):

not C

Mehek (mehek14):

and D so A, B, and D

OpenStudy (anonymous):

so i was right i was just missing one :)

Mehek (mehek14):

yup

OpenStudy (anonymous):

@pooja195 can you help me?

OpenStudy (anonymous):

or @Mehek14

Mehek (mehek14):

with?